# RESEARCH ARTICLE

### v1

@@ Abstract
[ABSTRACT]Epilepsy significantly impacts global health, affecting about 65 million people worldwide, along with various animal species. The diagnostic processes of epilepsy are often hindered by the transient and unpredictable nature of seizures. Here we propose a multi-space alignment approach based on cross-species and cross-modality electroencephalogram (EEG) data to enhance the detection capabilities and understanding of epileptic seizures. By employing deep learning techniques, including domain adaptation and knowledge distillation, our framework aligns cross-species and cross-modality EEG signals to enhance the detection capability beyond traditional within-species and with-modality models. Experiments on multiple surface and intracranial EEG datasets of humans and canines demonstrated substantial improvements in the detection accuracy, achieving over 90% AUC scores for cross-species and cross-modality seizure detection with extremely limited labeled data from the target species/modality. To our knowledge, this is the first study that demonstrates the effectiveness of integrating heterogeneous data from different species and modalities to improve EEG-based seizure detection performance. The approach may also be generalizable to different brain-computer interface paradigms, and suggests the possibility to combine data from different species/modalities to increase the amount of training data for large EEG models.
@@ 1 INTRODUCTION
Accurate decoding of epileptic EEG signals is crucial for aiding medical diagnosis, assisting neurologists in treatment, and reducing the risk of seizures[7]. Although EEG-based seizure detection has achieved significant progresses, these advances have imposed a considerable burden on physicians, requiring them to visually scrutinize up to several days of EEG signals to identify abnormal electrical discharges[9]. Such manual analysis is time-consuming and heavily expert-dependent, driving the recent development of automated algorithms for EEG-based epilepsy detection.
Previous studies involving traditional feature-based machine learning algorithms[10, 11, 12] and deep learning algorithms[13, 14, 15] have been widely explored for automatic epilepsy detection. Most works in the literature focus on within-patient seizure classification[13, 14, 15], in which the training and test sets typically originate from the same group of patients. However, such within-patient analysis could not be directly applied to unseen patients due to significant non-stationarity and individual differences of EEG signals[16]. Consequently, researchers have started investigating cross-patient seizure detection, utilizing auxiliary patient’s labeled data to relieve the reliance on target patient’s labeled data[16, 17, 18]. However, both within-patient and cross-patient seizure detection require an adequate amount of labeled data from the human patients for effective model training, which may not always be available.
@@
Structural and functional elements of biological systems are shared among vertebrates[20]. Canine epilepsy shares many characteristics with human epilepsy[23], including periodic features[3], energy features, and temporal morphology, as illustrated in Figures 1 a-c. Previous works have shown that human drugs such as levetiracetam and gabapentin are of benefit to dogs with refractory epilepsy[24], and naturally occurring canine epilepsy offers a potential biological model of human epilepsy[25]. Cross-species approaches to brain research can utilize prior knowledge from other species to enhance the understanding of diseases in the current species[20].
@@
Importantly, disparities exist in the EEG signals across species that must be appropriately addressed to achieve optimal transfer performance, which is the core problem we aim to solve in this work. EEG signals of different species are collected from varying numbers of electrodes, using different sampling rates and diverse collection devices, as illustrated in Figure 1d. These result in distinct signal characteristics. For instance, human iEEG signals could be sampled from various subdural electrode grids based on individual clinical considerations; the number of electrodes could go as many as 72, whereas canine iEEG signals may be acquired from an implanted device with 16 subdural electrodes[26]. In another example, the human iEEG signals were collected from three focal and three extra-focal electrodes[27], whereas sEEG signals were collected via scalp electrodes[28, 16]. These discrepancies complicate the cross-species transfer of feature heterogeneity. A naïve solution would be to discard the additional electrodes, resulting in the loss of valuable information. Even if the data are perfectly normalized in the input space, the extracted features may still exhibit significant distribution shifts, as illustrated in Figure 1e. Such problems have caused considerable difficulty in algorithm design for cross-species and cross-modality transfer.
To minimize cross-species and cross-modality discrepancies and enhance the transfer performance, we implemented alignments simultaneously in the input, feature and output spaces, eliminating the gaps of channel heterogeneity, distribution shifts, and prediction inconsistency. The details are illustrated in Figure  A1 of Appendix A.
Utilizing cross-species and cross-modality data could enhance the decoding of EEG signals for species with limited available data. Additionally, merging EEG datasets from different species and modalities can substantially increase the quantity of the training set, providing adequate training data for deep learning models or large-scale models. By exploring methods to mitigate data distribution discrepancies between species and modalities, we can also bridge inter-species and inter-modality gaps and offer novel insights for clinical endeavors.
@@ 2.3 Transfer learning approaches
We denote the proposed multi-space alignment approach as ResizeNet+MSA, which simultaneously performs input, feature, and output space alignments to handle channel disparity and distribution shifts for cross-species and cross-modality transfer.
@@ 2.4 Results
Our multi-space alignment framework is much more effective than all standalone strategies The multi-space alignment framework, encompassing simultaneously input, feature, and output space alignments, demonstrating superior performance over all standalone strategies. As shown in the ablation study results in Figure 2a, performing alignments in more spaces generally resulted in higher AUCs.
Figure 2b shows $t$-distributed stochastic neighbor embedding ($t$-SNE) visualizations[32] of iEEG features from canines and humans of the Kaggle dataset. In the two-dimensional reduced feature space, the distribution of raw iEEG data from different species exhibited clear distinctions (Baseline). However, following alignment by EA, domain adaptation, or knowledge distillation, the distributions became similar, with class separability significantly enhanced (different colors of the same shape merged).
@@
Cross-species and cross-modality feature heterogeneity can be accommodated by ResizeNet projection and knowledge distillation The integration of ResizeNet and knowledge distillation addresses the cross-species and cross-modality feature heterogeneity effectively. ResizeNet’s ability to project and transform features into a unified space, coupled with knowledge distillation’s ability in refining and transferring essential information, proved crucial. As shown in Tables 1-3, ResizeNet almost always outperformed Baseline approaches which brutally match the channels by truncation. ResizeNet ensured that the entire feature space pertinent to seizure detection was preserved and accurately interpreted across species, resulting in significant detection performance improvement. Our proposed ResizeNet+MSA almost always achieved the best performance.
To visualize the impact of the proposed approach, Figure 3a plots the iEEG/sEEG trials before and after input space alignment via the proposed ResizeNet in the time domain, and Figure 3b plots their approximate entropy. iEEG/sEEG signals before and after ResizeNet exhibit similar temporal structures and entropy features, demonstrating the effectiveness of ResizeNet in maintaining the key information while unifying signal forms across species.
@@
[b] Approach Canine (iEEG, Kaggle) to Human (iEEG, Kaggle) Human (iEEG, Kaggle) to Canine (iEEG, Kaggle) 5% 10% 15% 20% Avg. 5% 10% 15% 20% Avg. Baseline Within1 66.28±2.5 72.18±1.0 77.81±1.2 79.20±1.0 73.87 74.08±1.0 77.78±3.1 84.05±1.0 87.34±1.3 80.81 Comb.2 83.34±0.4 87.99±0.5 91.27±0.5 91.90±0.8 88.63 83.12±2.1 87.46±1.0 90.32±2.8 93.37±3.2 88.57 DAN 81.96±0.6 88.01±0.6 87.32±0.9 89.90±1.1 86.80 84.29±1.6 85.97±2.6 91.64±0.9 94.69±1.1 89.15 JAN 81.34±1.1 86.79±1.0 88.52±0.2 89.69±0.5 86.59 83.43±0.8 87.37±2.6 92.30±1.5 93.85±1.7 89.24 SHOT 85.41±1.6 86.97±1.4 89.13±0.9 90.90±0.6 88.10 82.50±1.5 87.14±1.0 90.70±0.5 91.94±0.3 88.07 DSAN 81.76±0.9 84.73±0.6 86.66±1.0 86.94±0.8 85.02 81.58±1.7 86.12±2.0 91.61±0.7 94.09±1.6 88.35 MCC 86.64±2.0 91.52±0.9 91.50±1.0 93.31±0.5 90.74 89.37±1.2 90.52±0.1 93.30±1.0 95.37±0.8 92.14 ResizeNet+ AT 89.29±0.6 92.68±0.5 94.09±0.5 95.88±0.3 92.99 86.16±1.2 91.05±0.9 94.22±0.5 97.34±0.7 92.19 NST 88.72±1.0 91.43±0.7 94.13±0.5 94.76±0.5 92.26 85.24±1.3 90.13±2.1 92.17±1.4 94.18±1.1 90.43 SP 90.31±0.2 93.08±0.7 94.52±0.7 95.64±0.0 93.39 86.61±1.1 91.22±0.6 94.79±1.5 96.84±0.6 92.37 RKD 88.55±0.6 91.52±0.6 92.49±0.0 94.19±0.5 91.69 85.97±0.5 90.02±1.8 92.22±1.8 92.86±1.7 90.27 PKT 89.26±0.7 92.63±0.7 94.10±0.4 96.12±0.6 93.03 85.81±0.5 90.38±1.7 95.12±1.2 97.96±0.3 92.32 CC 88.32±2.2 93.16±0.4 94.08±0.5 95.65±0.5 92.80 87.19±0.9 92.09±0.4 95.33±0.1 95.74±0.0 92.59 MSA (ours) 91.10±1.0 92.90±0.1 94.78±0.1 95.75±0.1 93.63 88.44±0.4 92.93±0.1 95.59±0.8 96.72±0.9 93.42 Approach Canine (iEEG, Kaggle) to Human (iEEG, Freiburg) Human (iEEG, Freiburg) to Canine (iEEG, Kaggle) 5% 10% 15% 20% Avg. 5% 10% 15% 20% Avg. Baseline Within 62.13±0.5 67.33±0.2 72.00±1.2 73.94±0.6 68.85 74.08±1.0 77.78±3.1 84.05±1.0 87.34±1.3 80.81 Comb. 83.86±2.5 84.62±0.3 85.84±0.8 85.61±0.6 84.98 87.03±0.4 92.98±0.4 95.47±0.7 96.15±0.3 92.91 DAN 89.81±1.3 92.45±0.9 93.16±1.2 93.44±0.7 92.22 85.26±1.7 89.42±1.7 93.68±0.4 93.73±0.5 90.52 JAN 86.47±1.3 89.65±0.3 89.28±0.1 90.81±0.5 89.05 85.71±0.3 90.12±0.2 94.57±0.4 94.79±0.6 91.30 SHOT 79.13±0.4 82.49±0.7 85.20±0.6 84.60±1.0 82.86 88.37±1.1 91.55±0.4 94.17±0.6 94.76±0.4 92.21 DSAN 84.87±3.0 86.42±1.5 86.22±2.0 88.70±1.4 86.55 80.52±0.6 85.84±1.1 89.95±1.6 88.19±2.8 86.13 MCC 85.69±0.8 86.59±1.0 86.96±0.5 87.79±0.8 86.76 90.69±0.6 95.93±0.4 96.50±0.3 97.29±0.7 95.10 ResizeNet+ AT 86.35±1.1 89.07±0.7 89.07±0.5 89.38±0.2 88.47 89.29±4.3 94.58±0.4 96.50±0.9 96.91±0.8 94.32 NST 84.55±1.7 86.38±1.1 87.32±2.1 87.95±2.3 86.55 88.17±4.1 88.88±3.3 93.74±1.9 93.78±2.7 91.14 SP 86.40±1.6 89.73±0.6 89.25±1.1 89.66±0.4 88.76 90.97±3.0 92.47±0.8 96.23±1.1 97.53±0.1 94.30 RKD 82.84±1.8 88.06±1.0 87.68±2.2 88.99±2.1 86.89 88.84±1.7 93.48±2.3 96.90±0.2 97.21±0.4 94.11 PKT 86.47±1.2 89.65±0.3 89.15±0.0 90.77±0.5 89.01 91.59±2.3 94.37±1.1 96.84±1.2 97.32±0.2 95.03 CC 86.12±0.6 89.32±0.6 89.57±0.2 90.43±0.7 88.86 89.53±3.9 92.95±2.0 96.33±0.2 96.94±0.6 93.94 MSA (ours) 91.06±0.9 93.90±1.2 94.78±1.0 94.79±0.8 93.63 92.85±0.5 96.92±0.9 97.04±0.6 97.92±0.1 96.18
@@
Table 3: Average semi-supervised cross-species transfer AUCs (%) on CHSZ and NICU sEEG datasets. The best average AUC of each task is marked in bold, and the second best by an underline.
@@
Figure 2: Ablation study, $t$-SNE feature visualization, and parameter sensitivity analysis. a, Ablation study on the Kaggle iEEG dataset, including two tasks of canine-to-human and human-to-canine, and two scenarios of unsupervised cross-species transfer and semi-supervised cross-species transfer. b, Feature $t$-SNE visualizations on the Kaggle iEEG data. c, Parameter sensitivity analysis of $\lambda$ and $\beta$ on the Kaggle iEEG dataset and CHSZ sEEG dataset using the proposed ResizeNet+MSA approach. A point denotes the average, and the shadow denotes the standard deviation.
Figure 3: Effect of input space alignment via the proposed ResizeNet on the four datasets. a, The iEEG/sEEG signals before and after ResizeNet exhibit similar temporal structures, demonstrating the effectiveness of ResizeNet. ResizeNet unifies the number of channels from different species, accommodating the input space discrepancies while preserving critical temporal structures and features. b, Mean approximate entropy of all channels before and after ResizeNet on the four datasets, showing that ResizeNet preserves important entropy features of the iEEG/sEEG signals.

### v2

@@ Abstract
[ABSTRACT]Epilepsy significantly impacts global health, affecting about 65 million people worldwide, along with various animal species. The diagnostic processes of epilepsy are often hindered by the transient and unpredictable nature of seizures. Here we propose a multi-space alignment approach based on cross-species and cross-modality electroencephalogram (EEG) data to enhance the detection capabilities and understanding of epileptic seizures. By employing deep learning techniques, including domain adaptation and knowledge distillation, our framework aligns cross-species and cross-modality EEG signals to enhance the detection capability beyond traditional within-species and within-modality models. Experiments on multiple surfaces and intracranial EEG datasets of humans and canines demonstrated substantial improvements in detection accuracy, achieving over 90% AUC scores for cross-species and cross-modality seizure detection with extremely limited labeled data from the target species/modality. To our knowledge, this is the first study that demonstrates the effectiveness of integrating heterogeneous data from different species and modalities to improve EEG-based seizure detection performance. This is a pilot study that provides insights into the challenges and potential of multi-species and multi-modality data integration, offering an effective solution for future work to collect huge EEG data to train large brain models.
@@ 1 INTRODUCTION
Accurate decoding of epileptic EEG signals is crucial for aiding medical diagnosis, assisting neurologists in treatment, and reducing the risk of seizures[7]. Although EEG-based seizure detection has achieved significant progresses, these advances have imposed a considerable burden on physicians, requiring them to visually scrutinize up to several days of EEG signals to identify abnormal electrical discharges[9]. Modern deep learning approaches highly rely on large hand-labeled datasets, which has been a significant bottleneck in healthcare applications of deep learning [10]. Detecting epilepsy for a new dataset frequently demands labor-intensive annotation efforts specific to that dataset. Numerous medical machine learning models have demonstrated inflexibility in adapting to shifting patient populations, evolving disease manifestations, variations in acquisition hardware, and so on, rendering them less effective in practice. Furthermore, ensuring the model to generalize well over diverse patient populations with varying characteristics poses a significant challenge, often necessitating model retraining [7]. Such time-consuming and expertise-dependent process underscores the need for automated algorithms capable of reliable and transferable EEG-based epilepsy detection.
Previous studies involving traditional feature-based machine learning algorithms[11, 12, 13] and deep learning algorithms[14, 15, 16] have been widely explored for automatic epilepsy detection. Most works in the literature focused on within-patient seizure classification[14, 15, 16], in which the training and test sets typically originate from the same group of patients. However, such within-patient analysis could not be directly applied to unseen patients due to significant non-stationarity and individual differences of EEG signals[17]. Consequently, researchers have started investigating cross-patient seizure detection, utilizing auxiliary patient’s labeled data to relieve the reliance on target patient’s labeled data[17, 18, 19]. However, both within-patient and cross-patient seizure detection require an adequate amount of labeled data from the human patients for effective model training, which may not always be available.
@@
Structural and functional elements of biological systems are shared among vertebrates[21]. Canine epilepsy shares many characteristics with human epilepsy[24], including seizure clusters [25], clinical presentation [26], pharmacological features [27], periodic features[3, 25], energy features, and temporal morphology [28], as illustrated in Figures 1 a-c. For periodic features, Gregg et al. [25] demonstrated that seizure periodicities and clusters are prevalent in canines, as in humans. For pharmacological features, Leppik et al. [27] observed that naturally occurring canine seizures closely mirror human epilepsy in pathophysiology, clinical progression, and response to conventional therapies. Additionally, canine studies on dosing and plasma drug concentrations for controlling seizures offer valuable insights for optimizing human treatment regimens. For brain wave characteristics, Pellegrino et al. [28] found that canine brain wave patterns closely resemble those of humans and apes, enabling cross-species comparisons to extrapolate healthy and seizure states. Furthermore, canine’s brain size makes them compatible with iEEG devices designed for human use [29]. Previous works have shown that human drugs such as levetiracetam and gabapentin are of benefit to canines with refractory epilepsy[30], and naturally occurring canine epilepsy offers a potential biological model of human epilepsy[31]. Adapted human and murine assays are effectively employed to examine the presence of neural antibodies in canines with idiopathic epilepsy [32]. Similarities are found in diagnostic biomarkers between the human and canine species [33]. Cross-species approaches to brain research can utilize prior knowledge from other species to enhance the understanding of diseases in the current species[21].
@@
Importantly, disparities exist in the EEG signals across species that must be appropriately addressed to achieve optimal transfer performance, which is the core problem we aim to solve in this work. EEG signals of different species are collected from varying numbers of electrodes, using different sampling rates and diverse collection devices, as illustrated in Figure 1d. These factors lead to distinct signal characteristics, manifesting as differences in brain signals’ temporal and spatial dimensions across different species and modalities. For instance, human iEEG signals could be sampled from various subdural electrode grids based on individual clinical considerations; the number of electrodes could go as many as 72, whereas canine iEEG signals may be acquired from an implanted device with 16 subdural electrodes[34]. In another example, the human iEEG signals were collected from three focal and three extra-focal electrodes[35], whereas sEEG signals were collected via scalp electrodes[36, 17]. These discrepancies complicate the cross-species transfer of feature heterogeneity. A naïve solution would be to discard the additional electrodes, resulting in the loss of valuable information. Even if the data are perfectly normalized in the input space, the extracted features may still exhibit significant distribution shifts, as illustrated in Figure 1e. Such problems have caused considerable difficulty in algorithm design for cross-species and cross-modality transfer.
In summary, the following technical challenges inherent in cross-species and cross-modality transfer must be systematically addressed:
  1. 1.
    Input heterogeneity across species and modalities: Differences in electrode configurations, sampling rates, and signal properties present significant obstacles to aligning the input spaces.
  2. 2.
    Distribution discrepancies across species, datasets, and subjects: Even after addressing the input heterogeneity, variations persist across input, feature, and output spaces. These differences pose additional challenges for effective transfer.
  3. 3.
    Limited labeled data in the target species, a common yet critical limitation in automatic seizure detection.
To minimize cross-species and cross-modality discrepancies and enhance the transfer performance, we implement alignments simultaneously in the input, feature and output spaces, eliminating the gaps of channel heterogeneity, distribution shifts, and prediction inconsistency. The details are illustrated in Figure  A1 of Appendix A.
This study is important, because animal brain signals, such as canine EEG data, provide valuable complementary insights due to cross-species similarities in seizure patterns and underlying neural mechanisms. Utilizing cross-species and cross-modality data could enhance the decoding of EEG signals for species with limited available data. Additionally, merging EEG datasets from different species and modalities can substantially increase the quantity of the training set, providing adequate training data for building robust and general-purpose deep learning models or large-scale models. By exploring methods to mitigate data distribution discrepancies between species and modalities, we can also bridge inter-species and inter-modality gaps and offer novel insights for clinical endeavors.
@@ 2.3 Transfer learning approaches
We denote the proposed multi-space alignment (MSA) approach as ResizeNet+MSA, which simultaneously performs input, feature, and output space alignments to handle channel disparity and distribution shifts for cross-species and cross-modality transfer.
@@ 2.4 Results
Our MSA framework is much more effective than all standalone strategies The MSA framework encompasses simultaneously input, feature, and output space alignments, demonstrating superior performance over all standalone strategies. As shown in the ablation study results in Figure 2a, performing alignments in more spaces generally resulted in higher AUCs. The basic baselines, “Source Only” in the unsupervised transfer scenario and “Within” or “Comb.” in the semi-supervised transfer scenario, consistently demonstrated the poorest performance across all tasks due to the absence of any alignment strategies. EA (input space alignment) and MCC (feature space alignment) were effective across most tasks. ResizeNet+KD and ResizeNet+MSA consistently achieved the second-best or best performance. While ResizeNet+KD focused on input space and output space alignments, ResizeNet+MSA extended this by performing alignment across the input, feature and output spaces, incorporating unsupervised domain adaptation beyond what ResizeNet+KD provides.
Figure 2b presents $t$-distributed stochastic neighbor embedding ($t$-SNE) [40] visualizations of iEEG/sEEG features from the cross-species transfer task of Canine-to-Human on the Kaggle, CHSZ, and NICU datasets. In the two-dimensional reduced feature space, raw iEEG data from different species showed distinct distributions (Baseline). However, after alignment using EA, domain adaptation, or the proposed ResizeNet+MSA, the distributions became more similar, with significantly improved class separability (indicated by the merging of different colors for the same shape). Among these methods, ResizeNet+MSA achieved the best results, showing superior class separability (the circles were clearly distinguished from the crosses) and optimal alignment performance (blue and red points of the same shape were closely aligned).
@@
Cross-species and cross-modality feature heterogeneity can be accommodated by ResizeNet projection and knowledge distillation The integration of ResizeNet and knowledge distillation addresses the cross-species and cross-modality feature heterogeneity effectively. ResizeNet’s ability to project and transform features into a unified space, coupled with knowledge distillation’s ability in refining and transferring essential information, proved crucial. As shown in Tables 1-3, ResizeNet almost always outperformed Baseline approaches which brutally match the channels by truncation. ResizeNet ensured that the entire feature space pertinent to seizure detection was preserved and accurately interpreted across species, resulting in significant detection performance improvement.
Our proposed ResizeNet+MSA always achieved the best average performance in Tables 1 and 2. Although in Table 3 the average performance of ResizeNet+MSA was worse than the six ResizeNet based approaches in transferring from Canine (iEEG, Kaggle) to Human (sEEG, CHSZ), its average performance over the four transfer scenarios in Table 3 was still the best among the 14 approaches (the average accuracy of ResizeNet+MSA was 87.45%, and the average accuracy of the second best approach, ResiezeNet+CC, was 86.27%). Moreover, ResizeNet+MSA was the only approach that consistently outperformed all Baselines, superior to all six other ResizeNet based approaches. These results demonstrated the effectiveness of MSA.
ResizeNet preserves essential classification information To assess the impact of the proposed ResizeNet, Figure 3a shows the iEEG/sEEG trials in the time domain before and after input space alignment, Figure 3b illustrates their approximate entropy, and Figure 3c shows the spatial characteristics of channel importance before and after applying ResizeNet.
The importance of each EEG channel was evaluated using SHapley Additive exPlanations (SHAP) values [41]. SHAP, an additive explanation model inspired by cooperative game theory, quantifies each feature’s (in this case, each channel’s) contribution to the model’s predictions. Unlike traditional approaches for evaluating feature importance, which only identify significant features, SHAP also provides the magnitude and direction (positive or negative influence) of each feature’s impact on individual predictions.
The channel importance calculation pipeline included three main steps. First, approximate entropy features were extracted from each channel, producing $C$ features for each signal, originally of dimensionality $(C,T)$. Next, a Random Forest classifier was trained on all available samples. Finally, SHAP values were calculated for each feature (channel), and the mean SHAP value across all samples was used to determine the importance of each channel. We utilized the TreeExplainer tool from [42] to generate the visualizations in Figure 3c. In the plot, each row corresponds to a feature representing a specific EEG channel. The horizontal axis shows the SHAP values, which quantify each channel’s contribution to the model’s predictions for individual samples. Each dot represents a sample, with its color reflecting the feature value (red for higher values and blue for lower values). Channels are ranked along the vertical axis in descending order of importance based on their relevance to the prediction outcomes.
The top row of Figure 3c presents SHAP-based channel importance for raw signals, and the bottom row shows the importance after applying ResizeNet for channel dimensionality alignment. For different subjects (different columns), the most important raw channel still exhibited high importance (among the top 3) after ResizeNet. For instance, channel 16 maintained high importance in the first column, channels 3 and 6 in the second column, channels 2 and 3 in the third column, and channels 15 and 16 in the fourth column. These patterns indicated that critical features were preserved post-resizing. Additionally, the SHAP value distribution remained largely consistent between the raw and resized signals. These results demonstrated that ResizeNet effectively retained the spatial characteristics of channel importance, ensuring robustness across various datasets and subjects.
In summary, Figure 3 shows that iEEG/sEEG signals before and after ResizeNet exhibit similar temporal structures, entropy features, and spatial characteristics, demonstrating the effectiveness of ResizeNet in preserving the essential information while unifying diverse brain signals across species.
@@
[b] Cross-Species Transfer Between Canine (Kaggle) and Human (Kaggle) Approach Canine (iEEG, Kaggle) to Human (iEEG, Kaggle) Human (iEEG, Kaggle) to Canine (iEEG, Kaggle) 5% 10% 15% 20% Avg. 5% 10% 15% 20% Avg. Baseline Within1 66.28±2.5 72.18±1.0 77.81±1.2 79.20±1.0 73.87 74.08±1.0 77.78±3.1 84.05±1.0 87.34±1.3 80.81 Comb.2 83.34±0.4 87.99±0.5 91.27±0.5 91.90±0.8 88.63 83.12±2.1 87.46±1.0 90.32±2.8 93.37±3.2 88.57 DAN 81.96±0.6 88.01±0.6 87.32±0.9 89.90±1.1 86.80 84.29±1.6 85.97±2.6 91.64±0.9 94.69±1.1 89.15 JAN 81.34±1.1 86.79±1.0 88.52±0.2 89.69±0.5 86.59 83.43±0.8 87.37±2.6 92.30±1.5 93.85±1.7 89.24 SHOT 85.41±1.6 86.97±1.4 89.13±0.9 90.90±0.6 88.10 82.50±1.5 87.14±1.0 90.70±0.5 91.94±0.3 88.07 DSAN 81.76±0.9 84.73±0.6 86.66±1.0 86.94±0.8 85.02 81.58±1.7 86.12±2.0 91.61±0.7 94.09±1.6 88.35 MCC 86.64±2.0 91.52±0.9 91.50±1.0 93.31±0.5 90.74 89.37±1.2 90.52±0.1 93.30±1.0 95.37±0.8 92.14 ResizeNet+ AT 89.29±0.6 92.68±0.5 94.09±0.5 95.88±0.3 92.99 86.16±1.2 91.05±0.9 94.22±0.5 97.34±0.7 92.19 NST 88.72±1.0 91.43±0.7 94.13±0.5 94.76±0.5 92.26 85.24±1.3 90.13±2.1 92.17±1.4 94.18±1.1 90.43 SP 90.31±0.2 93.08±0.7 94.52±0.7 95.64±0.0 93.39 86.61±1.1 91.22±0.6 94.79±1.5 96.84±0.6 92.37 RKD 88.55±0.6 91.52±0.6 92.49±0.0 94.19±0.5 91.69 85.97±0.5 90.02±1.8 92.22±1.8 92.86±1.7 90.27 PKT 89.26±0.7 92.63±0.7 94.10±0.4 96.12±0.6 93.03 85.81±0.5 90.38±1.7 95.12±1.2 97.96±0.3 92.32 CC 88.32±2.2 93.16±0.4 94.08±0.5 95.65±0.5 92.80 87.19±0.9 92.09±0.4 95.33±0.1 95.74±0.0 92.59 MSA (ours) 91.10±1.0 92.90±0.1 94.78±0.1 95.75±0.1 93.63 88.44±0.4 92.93±0.1 95.59±0.8 96.72±0.9 93.42 Cross-Species Transfer Between Canine (Kaggle) and Human (Freiburg) Approach Canine (iEEG, Kaggle) to Human (iEEG, Freiburg) Human (iEEG, Freiburg) to Canine (iEEG, Kaggle) 5% 10% 15% 20% Avg. 5% 10% 15% 20% Avg. Baseline Within 62.13±0.5 67.33±0.2 72.00±1.2 73.94±0.6 68.85 74.08±1.0 77.78±3.1 84.05±1.0 87.34±1.3 80.81 Comb. 83.86±2.5 84.62±0.3 85.84±0.8 85.61±0.6 84.98 87.03±0.4 92.98±0.4 95.47±0.7 96.15±0.3 92.91 DAN 89.81±1.3 92.45±0.9 93.16±1.2 93.44±0.7 92.22 85.26±1.7 89.42±1.7 93.68±0.4 93.73±0.5 90.52 JAN 86.47±1.3 89.65±0.3 89.28±0.1 90.81±0.5 89.05 85.71±0.3 90.12±0.2 94.57±0.4 94.79±0.6 91.30 SHOT 79.13±0.4 82.49±0.7 85.20±0.6 84.60±1.0 82.86 88.37±1.1 91.55±0.4 94.17±0.6 94.76±0.4 92.21 DSAN 84.87±3.0 86.42±1.5 86.22±2.0 88.70±1.4 86.55 80.52±0.6 85.84±1.1 89.95±1.6 88.19±2.8 86.13 MCC 85.69±0.8 86.59±1.0 86.96±0.5 87.79±0.8 86.76 90.69±0.6 95.93±0.4 96.50±0.3 97.29±0.7 95.10 ResizeNet+ AT 86.35±1.1 89.07±0.7 89.07±0.5 89.38±0.2 88.47 89.29±4.3 94.58±0.4 96.50±0.9 96.91±0.8 94.32 NST 84.55±1.7 86.38±1.1 87.32±2.1 87.95±2.3 86.55 88.17±4.1 88.88±3.3 93.74±1.9 93.78±2.7 91.14 SP 86.40±1.6 89.73±0.6 89.25±1.1 89.66±0.4 88.76 90.97±3.0 92.47±0.8 96.23±1.1 97.53±0.1 94.30 RKD 82.84±1.8 88.06±1.0 87.68±2.2 88.99±2.1 86.89 88.84±1.7 93.48±2.3 96.90±0.2 97.21±0.4 94.11 PKT 86.47±1.2 89.65±0.3 89.15±0.0 90.77±0.5 89.01 91.59±2.3 94.37±1.1 96.84±1.2 97.32±0.2 95.03 CC 86.12±0.6 89.32±0.6 89.57±0.2 90.43±0.7 88.86 89.53±3.9 92.95±2.0 96.33±0.2 96.94±0.6 93.94 MSA (ours) 91.06±0.9 93.90±1.2 94.78±1.0 94.79±0.8 93.63 92.85±0.5 96.92±0.9 97.04±0.6 97.92±0.1 96.18
@@
Table 3: Average semi-supervised cross-species and cross-modality transfer AUCs (%) on CHSZ and NICU sEEG datasets. The best average AUC of each task is marked in bold, and the second best by an underline.
Cross-Species and cross-modality Transfer Between Canine (Kaggle) and Human (CHSZ)
@@
Cross-Species and cross-modality Transfer Between Canine (Kaggle) and Human (NICU)
@@
Figure 2: Ablation study, $t$-SNE feature visualization, and parameter sensitivity analysis. a, Ablation study on the Kaggle iEEG dataset, including two tasks of canine-to-human and human-to-canine, and two scenarios of unsupervised cross-species transfer and semi-supervised cross-species transfer. b, Feature $t$-SNE visualizations on the Canine-to-Human cross-species transfer task of Kaggle, CHSZ and NICU datasets. c, Parameter sensitivity analysis of $\lambda$ and $\beta$ on the Kaggle iEEG dataset and CHSZ sEEG dataset using the proposed ResizeNet+MSA approach. A point denotes the average, and the shadow denotes the standard deviation.
Figure 3: Effect of Input Space Alignment via the Proposed ResizeNet on the Four Datasets (ResizeNet was not applied to the Kaggle H2 and H8 datasets, as their number of channels matches that of the canine dataset) a, The iEEG/sEEG signals before and after applying ResizeNet exhibit similar temporal structures, demonstrating its effectiveness. ResizeNet standardizes the number of channels across species, addressing input space discrepancies while preserving essential temporal structures and features. b, The mean approximate entropy of all channels before and after ResizeNet is shown for the four datasets, illustrating that ResizeNet retains critical entropy features of the iEEG/sEEG signals. c, Channel importance before and after applying ResizeNet, highlighting its ability to preserve significant spatial characteristics. Each column represents a different subject.